APC is about to launch its new Wings 'N Things fast food nationally. The research department is convinced that Wings 'N Things will be a great success and wants to introduce it immediately in all AFC outlets without advertising. The marketing department sees "things" differently and wants to unleash an intensive advertising campaign. The advertising campaign will cost $100,000 and if successful will produce $950,000 in revenue. If the campaign is unsuccessful (there is a 30% chance it won't be), the revenue is estimated at only $200,000. If no advertising is used, the revenue is estimated at $400,000 with probability .8 if customers are receptive and $200,000 with probability .2 if they are not.
(a) Draw the associated decision tree.
(b) What course of action should AFC follow in launching the new product
Answer #1

EMV o using advertisement = (0.3*950000+0.7*200000)-100000= 325000

EMV of not using advaertisement = 0.8*400000+0.2*200000 = 360000

Since the Expected value of not using advertisement is higher, the advertisement should not be used.

success 950000 0.3 0.7 200000 Failure Use advertising - 10 0000 use Ло 0.8 Yooooo advertising 0.2 -2оороо